A comittee of five persons is to be seleted andomly from a group of five men and 10 women. a) find the probability that the comittee consists of 2 men and 3 women b) find the probability that comittee consisting pf all the womens.
Found 2 solutions by stanbon, Susmitha:
Answer by stanbon(75887)

A comittee of five persons is to be seleted andomly from a group of five men and 10 women.
a) find the probability that the comittee consists of 2 men and 3 women
Number of ways to select 2 men: 5C2 = 10
Number of ways to select 3 women: 10C3 = 120
---
Number of ways to select 2 men and 3 women: 10*120 = 1200
---
Number of random groups of 5 from 15 = 3003
P(select 2 men and 3 women) = 1200/3003 = 0.4
----------------------------------------------------
b) find the probability that comittee consisting pf all the womens.
Number of ways to select 5 women of 10 = 252
P(select 5 women) = 252 = 3003
==============================
Cheers,
Stan H.
==================

Answer by Susmitha(21)

A comittee of 5 persons to be selected from 5 men and 10 women
total=15
a comittee of 5 persons is to be selected fom a group of 15 members=5/15=P(AUB)
probability of 2 men and 3 women
P(A^B)=P(A)+P(B)-P(AUB)
=2/5 + 3/10 -5/15
=12+9-10/30
=11/30
b) the probability that comittee consists at all women is 5/10 =1/2